OpenStudy (anonymous):

matrices 1 by 2 can be multiplyed with matrices 2 by 1?

hartnn (hartnn):

matrix \(\Large a \times b \) can be multiplied with matrix \(\Large b \times c\)

hartnn (hartnn):

the requirement is that columns of first matrix should be equal to rows of 2nd matrix

Directrix (directrix):

Hint attached. http://www.mathsisfun.com/algebra/matrix-multiplying.html

OpenStudy (anonymous):



hartnn (hartnn):

\(\Large M_{a\times b } \times M_{b\times c} = M_{a\times c}\)

OpenStudy (perl):

yes they can